### v1

# Structure-Accurate Medical Image Translation based on Dynamic Frequency Balance and Knowledge Guidance

Abstract
Multimodal medical images play a crucial role in the precise and comprehensive clinical diagnosis. Diffusion model is a powerful strategy to synthesize the required medical images. However, existing approaches still suffer from the problem of anatomical structure distortion due to the overfitting of high-frequency information and the weakening of low-frequency information. Thus, we propose a novel method based on dynamic frequency balance and knowledge guidance. Specifically, we first extract the low-frequency and high-frequency components by decomposing the critical features of the model using wavelet transform. Then, a dynamic frequency balance module is designed to adaptively adjust frequency for enhancing global low-frequency features and effective high-frequency details as well as suppressing high-frequency noise. To further overcome the challenges posed by the large differences between different medical modalities, we construct a knowledge-guided mechanism that fuses the prior clinical knowledge from a visual language model with visual features, to facilitate the generation of accurate anatomical structures. Experimental evaluations on multiple datasets show the proposed method achieves significant improvements in qualitative and quantitative assessments, verifying its effectiveness and superiority.
1 Introduction
Multimodal images are crucial for comprehensive evaluation in medical image analysis, which can compensate for the limitations of a single modality and achieve a precise diagnosis [1, 2, 3]. In medical image analysis, especially brain tumor diagnosis, computed tomography (CT) and various magnetic resonance (MR) imaging techniques, including T1-weighted, T2-weighted, fluid-attenuated inversion recovery (FLAIR) and diffusion-weighted imaging (DWI), are widely used. Different modalities can provide different contrasts and complementary information for the clinical diagnosis [4, 5, 6, 7, 8, 9]. However, in the clinical setting, multimodal image acquisition is often limited by factors such as scanning costs and safety, resulting in missing data for some modalities. The lack of these key modalities can affect the accuracy of clinical diagnosis and treatment. Therefore, how to generate the missing modality using the existing modality has become a key issue in the field of medical research [10, 11, 12].
Figure 1: Schematic diagram of our research problem and solution. During medical imaging, certain modalities may be missing due to scanning cost and safety issues. Generating missing modalities via dynamic frequency balance and knowledge guidance can help radiologists make more comprehensive diagnoses.
Recently, diffusion models [13, 14] have shown outstanding performance in the field of image-to-image translation [15, 4, 16, 17]. It constructs a progressive generation framework by defining the forward process and the reverse process through rigorous mathematical derivation, with a more stable training process and a finer generation quality than GANs (Generative Adversarial Networks) [18]. However, it has been found that existing algorithms based on diffusion models have the obvious limitation of overfitting the high-frequency information, which makes it difficult for the model to fully capture the low-frequency information of the images. Besides, the generated images are prone to suffer from anatomical structure distortion due to the large differences between different medical modalities [19, 20].
To further improve image generation quality, researchers begin to introduce the prior knowledge of vision language model (VLM) into the generation process [21, 22, 23]. The pre-trained VLM based on the billion-parameter scale provides rich prior knowledge for image translation, greatly reducing the learning difficulty while ensuring that the generated images have the correct anatomical structure. However, existing algorithms often blindly accept all prior knowledge, resulting in a failure to effectively fuse prior knowledge with the model, which has become a challenge in image translation [24].
To address above issues, as shown in Fig. 1, we propose a structure-accurate translation method called DFBK, which is based on Dynamic Frequency Balance and Knowledge guidance. Specifically, we first extract the low-frequency and high-frequency components by decomposing the critical features of the model using wavelet transform. Subsequently, we designed a dynamic frequency balance module to adaptively adjust frequency for enhancing global low-frequency features and effective high-frequency details while suppressing high-frequency noise, thereby facilitating the model to generate correct anatomical structures. To address the challenges posed by the large differences between different medical modalities, we propose a novel knowledge-guided mechanism. The mechanism fuse prior clinical knowledge with global low-frequency features using pre-trained VLM. For high-frequency features, we only retain the critical prior clinical knowledge while suppressing irrelevant information to avoid the over-enhancement of high-frequency features. This mechanism further achieves frequency balance while fusing prior knowledge and avoids the anatomical structure distortion caused by blindly accepting all prior knowledge.
The main contributions are summarized as follows:
  • Diffusion model-based medical image translation can trigger anatomical structure distortion due to inappropriate utilization of frequency information, and we propose a novel method DFBK to mitigate this problem.
  • We design a dynamic frequency balance module to enhance low-frequency and useful high-frequency components while suppressing high-frequency noise, thus adaptively and optimally leveraging frequency information.
  • We construct a knowledge-guided mechanism that fuses prior clinical knowledge with model features, which effectively improves the translation ability against large differences between different medical modalities.
  • Our method is evaluated on multiple medical image translation tasks, which exhibits more accurate anatomical structures and superior quantitative metrics scores.
2 Related Work
GAN-based approaches have been widely used in image translation [25, 26, 27]. For example, Pix2pix [28] achieves pixel-level synthesis of paired images through a simple network; CycleGAN [29] introduces cyclic consistency loss to construct cross-domain image mapping. RevGAN [30] is further optimised by designing an approximately reversible architecture to maintain constant memory complexity. Ea-GAN [31] enhances the synthesis of 3D MR images by integrating edge information. ResViT [22] combines a Transformer architecture and fine-tuning techniques to capture contextual features. MaskGAN [32] introduces a mask generator to control the generated region for accurate structure mapping. In order to solve the problem of unstable quality of GAN generated images, SynDiff [17] designed a loop-consistent architecture using a diffusion model to achieve bidirectional inter-modal translation. ResShift[33] designed a diffusion model for direct translation, which reduces the number of diffusion steps and improves the translation efficiency. However, these approaches usually focus excessively on high-frequency features, resulting in inaccurate low-frequency anatomical structures. Therefore, a method is needed to balance the focus on different frequency information.
With the rapid development of VLM, text-driven approaches are gradually applied to image translation [34, 35]. StyleCLIP [36] enables CLIP-guided feature manipulation by inverting the source image and obtaining its latent encoding. I2I-Galip [21], on the other hand, uses pre-trained VLM, eliminating the need to individually train the generator-discriminator pairs for each pair of source-target images, thus improving the conversion efficiency and enhancing the accuracy of low-frequency anatomical structures. However, existing algorithms are crude in fusing prior knowledge and are unable to fuse effectively with models. Therefore, there is a need to explore a novel mechanism to effectively fuse prior knowledge to improve model performance.
3 Methodology
3.1 Preliminary
Notice. We use lower-case letters $x_{0}$ and $y_{0}$ to denote the images of target modality and source modality, respectively. Moreover, we use $t$ to represent the timestep that the diffusion model is going through and $T$ to represent the total number of timesteps. We use the diffusion model based on ResShift and set $t\in\{1,2,3,4\}$. We use $\mathcal{N}(\cdot;\mu,\sigma^{2}I)$ to denote a variable that obeys a Gaussian distribution with mean $\mu$ and variance matrix $\sigma^{2}$.
Discrete Wavelet Transformation (DWT). DWT decomposes the image into low-frequency and high-frequency components. The low-frequency component represents the anatomical structure, while the high-frequency components represent the edges and contour details of the image. In this paper, we use Haar wavelet for DWT, which is widely used in applications [37, 38].
Given an image or a feature $z\in\mathbb{R}^{H\times W\times C}$, its low-frequency component $z^{ll}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ and high-frequency components $z^{lh}$, $z^{hl}$ and $z^{hh}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ can be decomposed by DWT,
$z^{ll},z^{lh},z^{hl},z^{hh}=DWT(z).$ (1)
DWT can not only effectively decompose frequency components, but also reduce the size of input images or features to reduce computational costs and speed up inference.
ResShift. ResShift uses a custom efficient diffusion process to achieve image translation. Specifically, given a pair of images $x_{0}$ and $y_{0}$, the forward process gradually adds noise to $x_{0}$ to obtain a Gaussian distribution with $y_{0}$ as the mean. This is represented as follows,
$q(x_{t}|x_{t-1},y_{0})=\mathcal{N}\left(x_{t};x_{t-1}+\alpha_{t}e_{0},\kappa^{2}\alpha_{t}I\right),$ (2)
where $e_{0}=y_{0}-x_{0}$, which represents the residual between the images of source modality and target modality. $\alpha_{t}=\eta_{t}-\eta_{t-1}$, where $\{\eta_{t}\}^{T}_{t=1}$ is a sequence that increases monotonically with timestep $t$ and satisfies $\eta_{1}\to 0$ and $\eta_{T}\to 1$. $\kappa$ is a parameter controlling the variance. $I$ is the identity matrix.
Then, we can get the reverse process,
$q(x_{t-1}|x_{t},y_{0})=\mathcal{N}\left(x_{t-1};\mu_{\theta}(x_{t},y_{0},t),\kappa^{2}\frac{\eta_{t-1}}{\eta_{t}}\alpha_{t}I\right),$ (3)
where $\mu_{\theta}(x_{t},y_{0},t)=\frac{\eta_{t-1}}{\eta_{t}}x_{t}+\frac{\alpha_{t}}{\eta_{t}}f_{\theta}(x_{t},y_{0},t)$ and $f_{\theta}$ is a neural network with parameter $\theta$ used to predict $x_{0}$.
To train the network $f_{\theta}$, we can obtain the objective function as follows,
$\mathop{\min}_{\theta}\sum_{t}{\frac{\alpha_{t}}{2\kappa^{2}\eta_{t}\eta_{t-1}}||f_{\theta}(x_{t},y_{0},t)-x_{0}||^{2}_{2}}.$ (4)
3.2 Medical Image Translation with DFBK
In this section we will introduce the detailed structure of DFBK. The framework is shown in Fig. 2. Firstly, we introduce how to build the dynamic frequency balance module. Then, we will introduce the principles and structure of the knowledge-guided mechanism. Finally, we combine them to obtain a network framework based on dynamic frequency balance and knowledge guidance.
Figure 2: Framework of the proposed method DFBK. In the lower right corner is the Dynamic Frequency Balance module, which achieves frequency balance by adaptively enhancing low-frequency features and effective high-frequency features while suppressing high-frequency noise. In the lower left corner is the Knowledge-Guided mechanism, which fuses the prior knowledge with global low-frequency features and retains only critical prior knowledge and suppresses irrelevant information when fusing with high-frequency features to avoid over-enhancing high-frequency features. Embedding is obtained from the text information using VLM’s text encoder. By using the dynamic frequency balance module and the knowledge-guided mechanism, we can generate target modal images with accurate anatomical structures.
Dynamic Frequency Balance. To address the fact that existing approaches overfit high-frequency information and weaken low-frequency information, resulting in anatomical structure distortion, we proposed a dynamic frequency balance (DFB) module. As shown in Fig. 2, we first use Eq. 1 to decompose the features. Then, we use the self-attention mechanism [39] to enhance the global low-frequency features,
$z^{L}_{s}=S(\frac{Q_{z^{ll}}K_{z^{ll}}^{T}}{\sqrt{d}})V_{z^{ll}},$ (5)
where $S(\cdot)$ represents the $softmax$ function and $Q_{z^{ll}}$, $K_{z^{ll}}$, $V_{z^{ll}}$ represent the query, key and value obtained from $z^{ll}$ through a linear layer, respectively. $z_{s}^{L}$ is the enhanced low-frequency features and $d$ is the scaling factor.
In order to achieve dynamic frequency balance, we use the differential attention [40] mechanism for high-frequency features,
$z^{H}_{s}=\left(S(\frac{Q_{z^{H}}^{1}{K_{z^{H}}^{1}}^{T}}{\sqrt{d}})-\lambda\cdot S(\frac{Q_{z^{H}}^{2}{K_{z^{H}}^{2}}^{T}}{\sqrt{d}})\right)V_{z^{H}},$ (6)
where $z^{H}\in\mathbb{R}^{H\times W\times 3C}$ is obtained by concatenating $z^{lh}$, $z^{hl}$ and $z^{hh}$ according to channel dimensions. $Q_{z^{H}}^{1}$, $Q_{z^{H}}^{2}$, $K_{z^{H}}^{1}$, $K_{z^{H}}^{2}$, and $V_{z^{H}}$ are all obtained from $z^{H}$ through a linear layer, where $Q$ and $K$ represent queries and keys, respectively, and $V$ represents the value. $\lambda$ is a learnable parameter. In order to synchronize the learning dynamics, we re-parameterize $\lambda$ as follows,
$\lambda=e^{(\lambda_{q}^{1}\cdot\lambda_{k}^{1})}-e^{(\lambda_{q}^{2}\cdot\lambda_{k}^{2})}+\lambda_{init},$ (7)
where $\lambda_{q}^{1}$, $\lambda_{q}^{2}$, $\lambda_{k}^{1}$, $\lambda_{k}^{2}$ are learnable parameters, and $\lambda_{init}$ is a constant. Differential attention can dynamically filter out and enhance effective high-frequency information while suppressing useless high-frequency noise.
Then, we use the Inverse Wavelet Transformation (IWT) to recombine the frequency features,
$z_{s}=IWT(z^{L}_{s},z^{H}_{s}),$ (8)
where $z_{s}$ is the recombination feature. The use of the dynamic frequency balance module can effectively strengthen the model’s attention to low-frequency features, while dynamically adjusting the model’s attention area for high-frequency features to enhance useful information and suppress high-frequency noise.
Knowledge Guidance. In order to effectively introduce VLM’s prior clinical knowledge while ensuring the balance of frequency information, we designed a novel knowledge-guided mechanism (KG) as shown in Fig. 2, which is similar to the dynamic frequency balance module. Specifically, we first fuse prior clinical knowledge and low-frequency features decomposed by Eq. 1,
$z^{L}_{k}=S(\frac{Q_{z^{ll}}K_{c}^{T}}{\sqrt{d}})V_{c}^{1},$ (9)
where $K_{c}$ and $V_{c}^{1}$ represent the key and value obtained from the text embedding $c$ through a linear layer, and the text embedding is obtained from the text through the BiomedCLIP’s [41] text encoder.
Then, in order to avoid the over-enhancement of high-frequency features and ensure the frequency balance, we use differential cross attention to fuse high-frequency features with prior clinical knowledge,
$z^{H}_{k}=\left(S(\frac{Q_{z^{H}}^{1}{K_{c}^{1}}^{T}}{\sqrt{d}})-\lambda\cdot S(\frac{Q_{z^{H}}^{2}{K_{c}^{2}}^{T}}{\sqrt{d}})\right)V_{c}^{2},$ (10)
where $K_{c}^{1}$, $K_{c}^{2}$ and $V_{c}^{2}$ are all obtained from the text embedding through a linear layer.
Finally, we can obtain the recombination feature $z_{k}$,
$z_{k}=IWT(z^{L}_{k},z^{H}_{k}).$ (11)
Overall Framework. We choose SwinUNet [42] as the base framework. Since the frequency balance in the backbone features is particularly important [19], we use the frequency balance module on backbone features before skipping connection. In addition, introducing prior knowledge in all layers can lead to fusion mismatch, and introducing it in bottom layers is more effective [43], so we decide to use the knowledge-guided mechanism in bottom layers. The details of the overall process are presented in Algorithm 1.
Algorithm 1 DFBK for medical image translation
0:  $z$, $c$, $d$, $\lambda_{q}^{1}$, $\lambda_{q}^{2}$, $\lambda_{k}^{1}$, $\lambda_{k}^{2}$, $\lambda_{init}$.
0:  The recombination feature $z_{s}$ and $z_{k}$.
1:  $z^{ll},z^{lh},z^{hl},z^{hh}\leftarrow DWT(z)$
2:  $z^{H}\leftarrow concat(z^{lh},z^{hl},z^{hh})$
3:  $\lambda\leftarrow e^{(\lambda_{q}^{1}\cdot\lambda_{k}^{1})}-e^{(\lambda_{q}^{2}\cdot\lambda_{k}^{2})}+\lambda_{init}$
4:  if use DFB then
5:     $Q_{z^{ll}},K_{z^{ll}},V_{z^{ll}}\leftarrow Linear(z^{ll})$
6:     $z^{L}_{s}\leftarrow S(\frac{Q_{z^{ll}}K_{z^{ll}}^{T}}{\sqrt{d}})V_{z^{ll}}$
7:     $Q_{z^{H}},K_{z^{H}},V_{z^{H}}\leftarrow Linear(z^{H})$
8:     $z^{H}_{s}\leftarrow(S(\frac{Q_{z^{H}}^{1}{K_{z^{H}}^{1}}^{T}}{\sqrt{d}})-\lambda\cdot S(\frac{Q_{z^{H}}^{2}{K_{z^{H}}^{2}}^{T}}{\sqrt{d}}))V_{z^{H}}$
9:     $z_{s}\leftarrow IWT(z^{L}_{s},z^{H}_{s})$triangleright Dynamic Frequency Balance
10:  else if use KG then
11:     $Q_{z^{ll}}\leftarrow Linear(z^{ll})$, $K_{c},V_{c}^{1}\leftarrow Linear(c)$
12:     $z^{L}_{k}\leftarrow S(\frac{Q_{z^{ll}}K_{c}^{T}}{\sqrt{d}})V_{c}^{1}$
13:     $Q_{z^{H}}^{1},Q_{z^{H}}^{2}\leftarrow Linear(z^{H})$, $K_{c}^{1},K_{c}^{2},V_{c}^{2}\leftarrow Linear(c)$
14:     $z^{H}_{k}\leftarrow(S(\frac{Q_{z^{H}}^{1}{K_{c}^{1}}^{T}}{\sqrt{d}})-\lambda\cdot S(\frac{Q_{z^{H}}^{2}{K_{c}^{2}}^{T}}{\sqrt{d}}))V_{c}^{2}$
15:     $z_{k}\leftarrow IWT(z^{L}_{k},z^{H}_{k})$triangleright Knowledge Guidance
16:  else
17:     $z_{s},z_{k}\leftarrow z$
18:  end if
19:  return  $z_{s}$ and $z_{k}$
4 Implementation Details
Figure 3: Comparison of approaches on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of T1 to T2 and T1 to FLAIR translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated image and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa (e.g., blue).
Figure 4: Comparison of approaches on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of FLAIR to DWI and FLAIR to T1 translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Figure 5: Comparison of approaches on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of CT to MR and MR to CT translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated image and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Table 1: Quantitative comparison of different approaches on BraTs and ISLES. The values in the table represent the average values for the entire test dataset. Bold represents the best result and underlined represents the suboptimal result.
BraTs (2023) ISLES (2015)
Dataset T1–>T2 T1–>FLAIR FLAIR–>DWI FLAIR–>T1
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
Ea-GAN 22.43 0.825 0.083 21.16 0.773 0.091 24.18 0.850 0.065 21.35 0.810 0.092
RevGAN 22.61 0.816 0.080 21.83 0.769 0.085 25.40 0.862 0.056 23.64 0.844 0.072
MaskGAN 21.40 0.809 0.092 20.41 0.735 0.101 25.24 0.856 0.057 23.15 0.834 0.077
ResViT 22.89 0.822 0.080 21.87 0.786 0.085 25.67 0.872 0.055 24.90 0.873 0.064
SynDiff 21.68 0.814 0.091 20.41 0.755 0.100 25.39 0.872 0.056 24.32 0.867 0.069
ResShift 23.47 0.843 0.075 22.20 0.792 0.082 25.61 0.877 0.055 25.15 0.876 0.062
DFBK (ours) 23.61 0.846 0.074 22.38 0.797 0.080 25.76 0.879 0.054 25.42 0.879 0.060
4.1 Experimental Datasets
We utilize three public datasets for the evaluation and validation of DFBK. We use the Brain Tumor Segmentation challenge 2023 (BraTs 2023) [44, 45, 5, 46] dataset to perform common translation tasks such as T1 to T2 and T1 to FLAIR. In addition, we also perform experiments on rare translation tasks such as FLAIR to DWI and FLAIR to T1 using Ischemic Stroke Lesion Segmentation challenge 2015 (ISLES 2015) [47, 48]. To further verify the generalization ability, we also perform experiments on complex translation tasks of CT and MR using Synthesizing computed tomography for Radiotherapy 2023 (SynthRAD 2023) [49].
For BraTs dataset, we select 40,032 images from 1,251 subjects in the training set with 32 images per subject for model training. Additionally, 7,008 images were chosen from 219 subjects in the validation set with 32 images per subject for generating the target modality and evalution. For ISLES dataset, we select 1,792 images from 28 subjects with 64 images per subject for model training and 2,304 images from 36 subjects with 64 images per subject for evalution. For SynthRAD dataset, we select 9,216 images from 144 subjects with 64 images per subject for model training and 2,304 images from 36 subjects with 64 images per subject for evaluation. The more information on above datasets and training details is available in Supplementary Material A.
4.2 Comparison Approaches
We select three classes of approaches for comparison. The first is GANs-based approaches, EaGAN, RevGAN and MaskGAN. The second is transformer-based approach, ResViT, which is a large model fine-tuning approach. The third is diffusion-based approaches, SynDiff and ResShift. The training parameters are set according to the parameters specified in the original paper, respectively. We use three evaluation metrics, peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and root mean squared error (RMSE), to quantitatively evaluate the generated results to verify the effectiveness of our method.
5 Experimental Results
5.1 Qualitative Results
Fig. 3 shows the results of various methods for medical image translation on BraTs. Ea-GAN introduces grating-like artifacts, and the anatomical structure in some areas is very blurred or even completely distorted as shown in Fig. 3 (a). RevGAN has unrecognizable anatomical structures and introduces additional lesion areas (Fig. 3 (b)). MaskGAN suffers from generating incorrect content and the synthesized images appearing overly blurred (Fig. 3 (c)). ResViT generates a relatively correct anatomical structure, but the brightness of some areas does not match the target modality (Fig. 3 (d)). The anatomical structure generated by SynDiff is incorrect, and the brightness do not match the target modality (Fig. 3 (e)). ResShift generates a generally correct anatomical structure, but the inconsistency of the brightness of certain areas with the target modality is still a noticeable problem (Fig. 3 (f)). In contrast, the images generated by our proposed method demonstrate significantly improved anatomical accuracy and more precise contrast representation. As shown in Figure 3 (g), our approach effectively preserves the structural integrity of anatomical features while maintaining the correct brightness distribution according to target modality. Furthermore, our method achieves the better performance across the quantitative evaluation metrics, underscoring its superiority in medical image translation.
Table 2: Quantitative comparison of different approaches on SynthRAD. The values in the table represent the average values for the entire test dataset. Bold represents the best result and underlined represents the suboptimal result.
Dataset SynthRAD (2023)
MR–>CT CT–>MR
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
Ea-GAN 24.75 0.852 0.059 20.39 0.679 0.098
RevGAN 26.24 0.872 0.051 22.94 0.798 0.073
MaskGAN 25.56 0.871 0.055 22.04 0.761 0.081
ResViT 26.92 0.887 0.047 22.88 0.801 0.075
SynDiff 26.05 0.877 0.052 20.73 0.705 0.094
ResShift 27.31 0.896 0.045 23.35 0.824 0.071
DFBK (ours) 27.51 0.897 0.044 23.54 0.828 0.069
Table 3: Ablation study on Dynamic Frequency Balance (DFB) and Knowledge Guidance (KG). “None” indicates the performances in the absence of DFB and KG.
Task FLAIR–>DWI FLAIR–>T1
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
None 25.61 0.877 0.055 25.15 0.876 0.062
+DFB 25.70 0.878 0.055 25.21 0.877 0.062
+KG 25.68 0.877 0.055 25.31 0.878 0.061
+DFB, KG 25.76 0.879 0.055 25.42 0.879 0.060
Fig. 4 shows the results of various methods for medical image translation on ISLES. Ea-GAN cannot generate the correct anatomical structure, and the content is completely distorted, as shown in Figure 4 (a). RevGAN introduces faint raster-like artifacts, and some areas are distorted (Figure 4 (b)). MaskGAN generates blurry content (Figure 4 (c)). The anatomical structure generated by ResViT (Figure 4 (d)) and SynDiff (Figure 4 (e)) exhibits noticeable incorrect, with significant distortions in shape and structure. Additionally, the brightness in certain regions is excessively dark, which fails to align with the target modality. Similarly, the anatomical structure generated by ResShift (Figure 4 (f)) suffers from partial distortions, where some regions exhibit structural inconsistencies. Moreover, the brightness distribution is uneven, with certain areas deviating from the expected intensity levels of the target modality. In contrast, as shown in Figure 4 (g), our method presents more accurate anatomical structures and achieves better metric values.
Fig. 5 shows the results of various methods for medical image translation on SynthRAD. The task of CT to MR translation in SynthRAD is very difficult. Since CT images do not provide information about the internal anatomical structure of the brain, most approaches cannot generate correct or relevant target modality anatomical structures. As shown in Fig. 5 (a), Ea-GAN produces MR images with raster artifacts and lacks clear internal anatomical structures, while its CT images show blurred cranial boundaries. RevGAN can generate relevant MR structures but fails to delineate the lesion area, and its CT images also have blurred cranial boundaries (Fig. 5 (b)). MaskGAN confuses the lesion with normal brain structures in the MR images, and its CT images exhibit incorrect cranial boundaries (Fig. 5 (c)). Although ResViT (Fig. 5 (d)) and SynDiff (Fig. 5 (e)) localize the lesion area in generated MR images, they do not accurately reproduce the overall internal anatomy and CT images generated by ResViT miss part of the cranial boundaries. ResShift produces MR images with generally correct anatomy but still fails to locate the lesion area, and its CT images are also incomplete (Fig. 5 (f)). In contrast, our method accurately localizes the MR lesion area, generates correct anatomical structures, and produces CT images with correct cranial boundaries (Fig. 5 (g)). The more visualization results are in Supplementary Material B.
5.2 Quantitative Results
Tab. 1 shows the mean values for each quantitative metric, where higher PSNR and SSIM values perform better, while lower RMSE value performs better. Our method outperforms all other approaches in six translation tasks on three datasets. Its superior performance stems from effectively balancing the model’s extraction of high- and low-frequency information and effective prior knowledge guidance, thereby preventing anatomical distortions caused by too fast high-frequency convergence and insufficient low-frequency extraction. In contrast, GAN-based approaches produce poor results due to their inherent instability in training, which makes them particularly susceptible to outliers in low-quality datasets such as BraTs. SynDiff, which incorporates GAN principles into a diffusion model, also shows unsatisfactory performance because its more complex training process is more sensitive to data quality. ResViT achieves relatively good results by leveraging large model fine-tuning and incorporating prior knowledge for cross-domain image synthesis, and ResShift shows suboptimal performance on most metrics through its efficient diffusion process, but the images generated by ResShift still show brightness inconsistency and anatomical distortion due to insufficient low-frequency information extraction. In contrast, our method effectively achieves frequency balance and has better performance.
5.3 Ablation Study
In this section, we validate the effectiveness of dynamic frequency balance module and knowledge-guided mechanism. All experiments are performed on ISLES dataset.
Dynamic Frequency Balance module (DFB) and Knowledge-Guided mechanism (KG). We conduct ablation experiments on the dynamic frequency balance module and the knowledge-guided mechanism to verify their important roles in medical image translation.
From Tab. 3, we can see that both dynamic frequency balance and knowledge guidance play an important role in generating the correct anatomical structure and affect most indicators, such as PSNR and SSIM. At the same time, it can be seen that when the source modality and target modality are significantly different, such as the conversion from FLAIR to T1, the knowledge-guided mechanism plays an important role and can effectively guide the model to generate the correct anatomical structure.
Table 4: Ablation study on the components in the Dynamic Frequency Balance module. “None” indicates the performances when there are no additional operations in this module.
Task FLAIR–>DWI FLAIR–>T1
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
None 25.61 0.877 0.055 25.15 0.876 0.062
+L 25.65 0.878 0.055 25.17 0.876 0,063
+H 25.67 0.877 0.055 25.12 0.874 0,063
+NormAttn 25.59 0.878 0.055 25.09 0.874 0.063
+L, H 25.70 0.878 0.055 25.21 0.877 0.062
Table 5: Ablation study on the components in the Knowledge-Guided mechanism. “None” indicates the performances when there are no additional operations in this mechanism.
Task FLAIR–>DWI FLAIR–>T1
Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
None 25.61 0.877 0.055 25.15 0.876 0.062
+L 25.66 0.876 0.055 25.21 0.878 0,062
+H 25.59 0.875 0.055 25.15 0.874 0.062
+CrossAttn 25.64 0.877 0.055 25.21 0.875 0.062
+L, H 25.68 0.877 0.055 25.31 0.878 0.061
The components in DFB. In this section, we will verify the impact of low-frequency and high-frequency modules on model performance respectively.
It can be seen from Tab. 4 that only enhancing the low-frequency features (L) can alleviate the poor performance of the model due to too fast high-frequency convergence to a certain extent, while only dynamically balancing the high-frequency features (H) suppresses some high-frequency noise, but the enhancement of the effective high-frequency components actually accelerates the high-frequency convergence of the model and reduces the performance. The improvement of PSNR in the FLAIR to DWI task is related to its similar modality. To further demonstrate the importance of frequency balance, we replace the dynamic frequency balance module with a normal self-attention mechanism (NormAttn). The results show that the performance of the model has slightly decreased. This is because the self-attention mechanism can only enhance global features, but cannot balance high-frequency and low-frequency features, which makes the model further strengthen the convergence of high-frequency features. In contrast, the dynamic frequency balance module can decompose and process frequency components more carefully.
The components in KG. In order to verify the impact of each component of the knowledge-guided mechanism, we conducted an ablation experiment.
As shown in Tab. 5, only fusing the prior knowledge with low-frequency features can guide the model to enhance the extraction of low-frequency features and slightly improve the model performance. However, only fusing the prior knowledge with high-frequency features cannot achieve frequency balance, even if the enhancement of high-frequency features is suppressed. This shows that the enhancement of low-frequency is more important. To further verify the effectiveness of the knowledge-guided mechanism, we replaced it with a cross-attention mechanism (CrossAttn). The results show that the cross-attention mechanism can slightly improve the model performance, but it limits the guidance of the model by the prior knowledge due to the over-enhancement of high-frequency. In contrast, the knowledge-guided mechanism processes high and low frequencies separately, effectively improving model performance.
6 Conclusion
In this paper, we propose a new diffusion model architecture for medical image translation, which achieves generating target modalities with correct anatomical structures through the dynamic frequency balance module and the knowledge-guided mechanism. Experimental results demonstrate our method’s effectiveness in medical image translation and excellent performance in generating correct anatomical structures. Ablation experiments also demonstrate the effectiveness of our proposed dynamic frequency balance module and knowledge-guided mechanism.
Limitation. A major limitation of this study is that our method relies on high-quality paired data, but it may be difficult to obtain a large number of paired images due to the variety of factors, which may influence its broad applicability in some clinical scenarios. To overcome this limitation, our future research will focus on exploring medical image translation based on unpaired images.
In summary, we propose a novel method for structure-accurate medical image translation, which aims to overcome the interference of modality absence on clinical diagnosis. The proposed method is expected to facilitate radiologists in making more precise diagnoses.
Supplementary Material
Appendix A Experimental Details
A.1 Datasets
Brain Tumor Segmentation challenge 2023 (BraTs 2023). BraTs dataset has 1251 subjects for training and 219 subjects for evaluation, each subject has 155 slices and each slice size is 240 $\times$ 240. We select 40,032 images from 1,251 subjects with 32 images per subject in the training set for model training. Additionally, 7008 images were chosen from 219 subjects in the validation set, with 32 images per subject for generating the target modality and evalution. All images are center-cropped and resize to the size of 192 $\times$ 192 to remove redundant areas that do not contain brain information.
Ischemic Stroke Lesion Segmentation challenge 2015 (ISLES 2015). ISLES dataset has 28 subjects for training and 36 subjects for evaluation. The number and size of slices in each subject are inconsistent. We select 1,792 images from 28 subjects with 64 images per subject in the training set for model training. Additionally, 2,304 images were chosen from 36 subjects in the validation set, with 64 images per subject for generating the target modality and evalution. All images are center-cropped and resize to the size of 224 $\times$ 224 to remove redundant areas that do not contain brain information.
Synthesizing computed tomography for Radiotherapy 2023 (SynthRAD 2023). ISLES dataset has 144 subjects for training and 36 subjects for evaluation. The number and size of slices in each subject are inconsistent. We select 9,216 images from 144 subjects with 64 images per subject in the training set for model training, and 2,304 images were chosen from 36 subjects in the validation set, with 64 images per subject for generating the target modality and evalution. All images are center-cropped and resize to the size of 192 $\times$ 192 to remove redundant areas that do not contain brain information.
A.2 Training Details
Our method’s training is based on ResShift and SwinUNet. We set the number of the channel dimensions to 160, the number of head channels to 32, the number of residual block to 2, the swin depth to 2, and the swin embedding dim to 256. The channel multipliers are set to $\{1,1,2,2,4\}$ and $\{1,1,2,2,4,4\}$ for size 224 and 192, respectively. We set the batch size to 1 for ISLES dataset and 4 for BraTs and SynthRAD datasets, the total number of timesteps to 4, the window size to 14 for ISLES dataset and 8 for BraTs and SynthRAD datasets, and learning rate to $5.0\times 10^{-5}$ which is reduced by cosine to $2.0\times 10^{-5}$. We trained the model with 100k iterations. The diffusion model training is conducted on two RTX 4090 GPUs.
Appendix B Additional Experimental Results
As shown in Fig. 6, 7 and 8, our method consistently outperforms other methods for medical image translation on BraTs, ISLES, and SynthRAD datasets, demonstrating superior anatomical accuracy, and better quantitative metrics. Unlike other methods, our approach effectively maintains structural integrity, accurately locates lesion areas, and ensures realistic conversion between imaging modalities. Therefore, our method is robust and reliable in medical image translation and has great clinical application value.
Figure 6: Additional results on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of T1 to T2 and T1 to FLAIR translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Figure 7: Additional results on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of FLAIR to DWI and FLAIR to T1 translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
Figure 8: Additional results on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours). The first and third rows show the results of CT to MR and MR to CT translation, respectively. The first column is the source image, and the second column is the target ground-truth. The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively. PSNR and SSIM values of each image are shown in the corner of images. The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth. The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.

### v2

@@
-Structure-Accurate Medical Image Translation based on Dynamic Frequency Balance and Knowledge Guidance
+Structure-Accurate Medical Image Translation via Dynamic Frequency Balance and Knowledge Guidance
 Abstract
@@
 1 Introduction
+Figure 1: Schematic diagram of our research problem and solution.
+During medical imaging, certain modalities may be missing due to scanning cost and safety issues.
+Generating missing modalities via dynamic frequency balance and knowledge guidance can help radiologists make more comprehensive diagnoses.
 Multimodal images are crucial for comprehensive evaluation in medical image analysis, which can compensate for the limitations of a single modality and achieve a precise diagnosis [1, 2, 3].
@@
 Therefore, how to generate the missing modality using the existing modality has become a key issue in the field of medical research [10, 11, 12].
-Figure 1: Schematic diagram of our research problem and solution.
-During medical imaging, certain modalities may be missing due to scanning cost and safety issues.
-Generating missing modalities via dynamic frequency balance and knowledge guidance can help radiologists make more comprehensive diagnoses.
 Recently, diffusion models [13, 14] have shown outstanding performance in the field of image-to-image translation [15, 4, 16, 17].
@@
 To further improve image generation quality, researchers begin to introduce the prior knowledge of vision language model (VLM) into the generation process [21, 22, 23].
-The pre-trained VLM based on the billion-parameter scale provides rich prior knowledge for image translation, greatly reducing the learning difficulty while ensuring that the generated images have the correct anatomical structure.
+The pre-trained VLM provides rich prior knowledge for image translation, greatly reducing the learning difficulty while ensuring that the generated images have the correct anatomical structure.
 However, existing algorithms often blindly accept all prior knowledge, resulting in a failure to effectively fuse prior knowledge with the model, which has become a challenge in image translation [24].
@@
 We use $\mathcal{N}(\cdot;\mu,\sigma^{2}I)$ to denote a variable that obeys a Gaussian distribution with mean $\mu$ and variance matrix $\sigma^{2}$.
+Figure 2: Framework of the proposed method DFBK.
+In the lower right corner is the Dynamic Frequency Balance module, which achieves frequency balance by adaptively enhancing low-frequency features and effective high-frequency features while suppressing high-frequency noise.
+In the lower left corner is the Knowledge-Guided mechanism, which fuses the prior knowledge with global low-frequency features and retains only critical prior knowledge and suppresses irrelevant information when fusing with high-frequency features to avoid over-enhancing high-frequency features.
+Embedding is obtained from the text information using VLM’s text encoder.
+By using the dynamic frequency balance module and the knowledge-guided mechanism, we can generate target modal images with accurate anatomical structures.
 Discrete Wavelet Transformation (DWT).
@@
 3.2 Medical Image Translation with DFBK
-In this section we will introduce the detailed structure of DFBK.
-The framework is shown in Fig.
-2.
-Firstly, we introduce how to build the dynamic frequency balance module.
-Then, we will introduce the principles and structure of the knowledge-guided mechanism.
-Finally, we combine them to obtain a network framework based on dynamic frequency balance and knowledge guidance.
-Figure 2: Framework of the proposed method DFBK.
-In the lower right corner is the Dynamic Frequency Balance module, which achieves frequency balance by adaptively enhancing low-frequency features and effective high-frequency features while suppressing high-frequency noise.
-In the lower left corner is the Knowledge-Guided mechanism, which fuses the prior knowledge with global low-frequency features and retains only critical prior knowledge and suppresses irrelevant information when fusing with high-frequency features to avoid over-enhancing high-frequency features.
-Embedding is obtained from the text information using VLM’s text encoder.
-By using the dynamic frequency balance module and the knowledge-guided mechanism, we can generate target modal images with accurate anatomical structures.
-Dynamic Frequency Balance.
-To address the fact that existing approaches overfit high-frequency information and weaken low-frequency information, resulting in anatomical structure distortion, we proposed a dynamic frequency balance (DFB) module.
-As shown in Fig.
-2, we first use Eq.
-1 to decompose the features.
-Then, we use the self-attention mechanism [39] to enhance the global low-frequency features,
-$z^{L}_{s}=S(\frac{Q_{z^{ll}}K_{z^{ll}}^{T}}{\sqrt{d}})V_{z^{ll}},$ (5)
-where $S(\cdot)$ represents the $softmax$ function and $Q_{z^{ll}}$, $K_{z^{ll}}$, $V_{z^{ll}}$ represent the query, key and value obtained from $z^{ll}$ through a linear layer, respectively.
-$z_{s}^{L}$ is the enhanced low-frequency features and $d$ is the scaling factor.
-In order to achieve dynamic frequency balance, we use the differential attention [40] mechanism for high-frequency features,
-$z^{H}_{s}=\left(S(\frac{Q_{z^{H}}^{1}{K_{z^{H}}^{1}}^{T}}{\sqrt{d}})-\lambda\cdot S(\frac{Q_{z^{H}}^{2}{K_{z^{H}}^{2}}^{T}}{\sqrt{d}})\right)V_{z^{H}},$ (6)
-where $z^{H}\in\mathbb{R}^{H\times W\times 3C}$ is obtained by concatenating $z^{lh}$, $z^{hl}$ and $z^{hh}$ according to channel dimensions.
-$Q_{z^{H}}^{1}$, $Q_{z^{H}}^{2}$, $K_{z^{H}}^{1}$, $K_{z^{H}}^{2}$, and $V_{z^{H}}$ are all obtained from $z^{H}$ through a linear layer, where $Q$ and $K$ represent queries and keys, respectively, and $V$ represents the value.
-$\lambda$ is a learnable parameter.
-In order to synchronize the learning dynamics, we re-parameterize $\lambda$ as follows,
-$\lambda=e^{(\lambda_{q}^{1}\cdot\lambda_{k}^{1})}-e^{(\lambda_{q}^{2}\cdot\lambda_{k}^{2})}+\lambda_{init},$ (7)
-where $\lambda_{q}^{1}$, $\lambda_{q}^{2}$, $\lambda_{k}^{1}$, $\lambda_{k}^{2}$ are learnable parameters, and $\lambda_{init}$ is a constant.
-Differential attention can dynamically filter out and enhance effective high-frequency information while suppressing useless high-frequency noise.
-Then, we use the Inverse Wavelet Transformation (IWT) to recombine the frequency features,
-$z_{s}=IWT(z^{L}_{s},z^{H}_{s}),$ (8)
-where $z_{s}$ is the recombination feature.
-The use of the dynamic frequency balance module can effectively strengthen the model’s attention to low-frequency features, while dynamically adjusting the model’s attention area for high-frequency features to enhance useful information and suppress high-frequency noise.
-Knowledge Guidance.
-In order to effectively introduce VLM’s prior clinical knowledge while ensuring the balance of frequency information, we designed a novel knowledge-guided mechanism (KG) as shown in Fig.
-2, which is similar to the dynamic frequency balance module.
-Specifically, we first fuse prior clinical knowledge and low-frequency features decomposed by Eq.
-1,
-$z^{L}_{k}=S(\frac{Q_{z^{ll}}K_{c}^{T}}{\sqrt{d}})V_{c}^{1},$ (9)
-where $K_{c}$ and $V_{c}^{1}$ represent the key and value obtained from the text embedding $c$ through a linear layer, and the text embedding is obtained from the text through the BiomedCLIP’s [41] text encoder.
-Then, in order to avoid the over-enhancement of high-frequency features and ensure the frequency balance, we use differential cross attention to fuse high-frequency features with prior clinical knowledge,
-$z^{H}_{k}=\left(S(\frac{Q_{z^{H}}^{1}{K_{c}^{1}}^{T}}{\sqrt{d}})-\lambda\cdot S(\frac{Q_{z^{H}}^{2}{K_{c}^{2}}^{T}}{\sqrt{d}})\right)V_{c}^{2},$ (10)
-where $K_{c}^{1}$, $K_{c}^{2}$ and $V_{c}^{2}$ are all obtained from the text embedding through a linear layer.
-Finally, we can obtain the recombination feature $z_{k}$,
-$z_{k}=IWT(z^{L}_{k},z^{H}_{k}).$ (11)
-Overall Framework.
-We choose SwinUNet [42] as the base framework.
-Since the frequency balance in the backbone features is particularly important [19], we use the frequency balance module on backbone features before skipping connection.
-In addition, introducing prior knowledge in all layers can lead to fusion mismatch, and introducing it in bottom layers is more effective [43], so we decide to use the knowledge-guided mechanism in bottom layers.
-The details of the overall process are presented in Algorithm 1.
 Algorithm 1 DFBK for medical image translation
@@
 19:  return  $z_{s}$ and $z_{k}$
-4 Implementation Details
-Figure 3: Comparison of approaches on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours).
+In this section we will introduce the detailed structure of DFBK.
+The framework is shown in Fig.
+2.
+Firstly, we introduce how to build the dynamic frequency balance module.
+Then, we will introduce the principles and structure of the knowledge-guided mechanism.
+Finally, we combine them to obtain a network framework based on dynamic frequency balance and knowledge guidance.
+Dynamic Frequency Balance.
+To address the fact that existing approaches overfit high-frequency information and weaken low-frequency information, resulting in anatomical structure distortion, we proposed a dynamic frequency balance (DFB) module.
+As shown in Fig.
+2, we first use Eq.
+1 to decompose the features.
+Then, we use the self-attention mechanism [39] to enhance the global low-frequency features,
+$z^{L}_{s}=S(\frac{Q_{z^{ll}}K_{z^{ll}}^{T}}{\sqrt{d}})V_{z^{ll}},$ (5)
+where $S(\cdot)$ represents the $softmax$ function and $Q_{z^{ll}}$, $K_{z^{ll}}$, $V_{z^{ll}}$ represent the query, key and value obtained from $z^{ll}$ through a linear layer, respectively.
+$z_{s}^{L}$ is the enhanced low-frequency features and $d$ is the scaling factor.
+In order to achieve dynamic frequency balance, we use the differential attention mechanism for high-frequency features,
+$z^{H}_{s}=\left(S(\frac{Q_{z^{H}}^{1}{K_{z^{H}}^{1}}^{T}}{\sqrt{d}})-\lambda\cdot S(\frac{Q_{z^{H}}^{2}{K_{z^{H}}^{2}}^{T}}{\sqrt{d}})\right)V_{z^{H}},$ (6)
+where $z^{H}\in\mathbb{R}^{H\times W\times 3C}$ is obtained by concatenating $z^{lh}$, $z^{hl}$ and $z^{hh}$ according to channel dimensions.
+$Q_{z^{H}}^{1}$, $Q_{z^{H}}^{2}$, $K_{z^{H}}^{1}$, $K_{z^{H}}^{2}$, and $V_{z^{H}}$ are all obtained from $z^{H}$ through a linear layer, where $Q$ and $K$ represent queries and keys, respectively, and $V$ represents the value.
+$\lambda$ is a learnable parameter.
+In order to synchronize the learning dynamics, we re-parameterize $\lambda$ as follows,
+$\lambda=e^{(\lambda_{q}^{1}\cdot\lambda_{k}^{1})}-e^{(\lambda_{q}^{2}\cdot\lambda_{k}^{2})}+\lambda_{init},$ (7)
+where $\lambda_{q}^{1}$, $\lambda_{q}^{2}$, $\lambda_{k}^{1}$, $\lambda_{k}^{2}$ are learnable parameters, and $\lambda_{init}$ is a constant.
+Differential attention [40] can dynamically filter out and enhance effective high-frequency information while suppressing useless high-frequency noise.
+Then, we use the Inverse Wavelet Transformation (IWT) to recombine the frequency features,
+$z_{s}=IWT(z^{L}_{s},z^{H}_{s}),$ (8)
+where $z_{s}$ is the recombination feature.
+The use of the dynamic frequency balance module can effectively strengthen the model’s attention to low-frequency features, while dynamically adjusting the model’s attention area for high-frequency features to enhance useful information and suppress high-frequency noise.
+Figure 3: Comparison of approaches on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) DFBK (ours).
 The first and third rows show the results of T1 to T2 and T1 to FLAIR translation, respectively.
@@
 The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa (e.g., blue).
-Figure 4: Comparison of approaches on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours).
+Figure 4: Comparison of approaches on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours).
 The first and third rows show the results of FLAIR to DWI and FLAIR to T1 translation, respectively.
@@
 The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively.
-PSNR and SSIM values of each image are shown in the corner of images.
-The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated images and the ground truth.
-The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
-Figure 5: Comparison of approaches on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours).
-The first and third rows show the results of CT to MR and MR to CT translation, respectively.
-The first column is the source image, and the second column is the target ground-truth.
-The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively.
-PSNR and SSIM values of each image are shown in the corner of images.
-The yellow box indicates the zoomed-in visualization area, and the blue box represents the difference heatmap between the generated image and the ground truth.
-The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
-Table 1: Quantitative comparison of different approaches on BraTs and ISLES.
-The values in the table represent the average values for the entire test dataset.
-Bold represents the best result and underlined represents the suboptimal result.
-BraTs (2023) ISLES (2015)
-Dataset T1–>T2 T1–>FLAIR FLAIR–>DWI FLAIR–>T1
-Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
-Ea-GAN 22.43 0.825 0.083 21.16 0.773 0.091 24.18 0.850 0.065 21.35 0.810 0.092
-RevGAN 22.61 0.816 0.080 21.83 0.769 0.085 25.40 0.862 0.056 23.64 0.844 0.072
-MaskGAN 21.40 0.809 0.092 20.41 0.735 0.101 25.24 0.856 0.057 23.15 0.834 0.077
-ResViT 22.89 0.822 0.080 21.87 0.786 0.085 25.67 0.872 0.055 24.90 0.873 0.064
-SynDiff 21.68 0.814 0.091 20.41 0.755 0.100 25.39 0.872 0.056 24.32 0.867 0.069
-ResShift 23.47 0.843 0.075 22.20 0.792 0.082 25.61 0.877 0.055 25.15 0.876 0.062
-DFBK (ours) 23.61 0.846 0.074 22.38 0.797 0.080 25.76 0.879 0.054 25.42 0.879 0.060
+Knowledge Guidance.
+In order to effectively introduce VLM’s prior clinical knowledge while ensuring the balance of frequency information, we designed a novel knowledge-guided mechanism (KG) as shown in Fig.
+2, which is similar to the dynamic frequency balance module.
+Specifically, we first fuse prior clinical knowledge and low-frequency features decomposed by Eq.
+1,
+$z^{L}_{k}=S(\frac{Q_{z^{ll}}K_{c}^{T}}{\sqrt{d}})V_{c}^{1},$ (9)
+where $K_{c}$ and $V_{c}^{1}$ represent the key and value obtained from the text embedding $c$ through a linear layer, and the text embedding is obtained from the text through the BiomedCLIP’s [41] text encoder.
+Then, in order to avoid the over-enhancement of high-frequency features and ensure the frequency balance, we use differential cross attention to fuse high-frequency features with prior clinical knowledge,
+$z^{H}_{k}=\left(S(\frac{Q_{z^{H}}^{1}{K_{c}^{1}}^{T}}{\sqrt{d}})-\lambda\cdot S(\frac{Q_{z^{H}}^{2}{K_{c}^{2}}^{T}}{\sqrt{d}})\right)V_{c}^{2},$ (10)
+where $K_{c}^{1}$, $K_{c}^{2}$ and $V_{c}^{2}$ are all obtained from the text embedding through a linear layer.
+Finally, we can obtain the recombination feature $z_{k}$,
+$z_{k}=IWT(z^{L}_{k},z^{H}_{k}).$ (11)
+Through the dynamic frequency balance module and the knowledge guidance mechanism, we can balance the model’s ability to extract features of high and low frequency information, thereby generating accurate anatomical structures.
+Overall Framework.
+We choose SwinUNet [42] as the base framework.
+Since the frequency balance in the backbone features is particularly important [19], we use the frequency balance module on backbone features before skipping connection.
+In addition, introducing prior knowledge in all layers can lead to fusion mismatch, and introducing it in bottom layers is more effective [43], so we decide to use the knowledge-guided mechanism in bottom layers.
+The details of the overall process are presented in Algorithm 1.
+4 Implementation Details
 4.1 Experimental Datasets
@@
 To further verify the generalization ability, we also perform experiments on complex translation tasks of CT and MR using Synthesizing computed tomography for Radiotherapy 2023 (SynthRAD 2023) [49].
-For BraTs dataset, we select 40,032 images from 1,251 subjects in the training set with 32 images per subject for model training.
-Additionally, 7,008 images were chosen from 219 subjects in the validation set with 32 images per subject for generating the target modality and evalution.
-For ISLES dataset, we select 1,792 images from 28 subjects with 64 images per subject for model training and 2,304 images from 36 subjects with 64 images per subject for evalution.
-For SynthRAD dataset, we select 9,216 images from 144 subjects with 64 images per subject for model training and 2,304 images from 36 subjects with 64 images per subject for evaluation.
-The more information on above datasets and training details is available in Supplementary Material A.
+For BraTs dataset, we select 40,032 images from 1,251 subjects with 32 images per subject for training.
+Additionally, 7,008 images were chosen from 219 subjects with 32 images per subject for generating the target modality and evalution.
+For ISLES dataset, we select 1,792 images from 28 subjects with 64 images per subject for training and 2,304 images from 36 subjects with 64 images per subject for evalution.
+For SynthRAD dataset, we select 9,216 images from 144 subjects with 64 images per subject for training and 2,304 images from 36 subjects with 64 images per subject for evaluation.
+The more information is available in Supplementary Material A.
 4.2 Comparison Approaches
@@
 The second is transformer-based approach, ResViT, which is a large model fine-tuning approach.
-The third is diffusion-based approaches, SynDiff and ResShift.
+The third is SynDiff whicth is based on diffusion model.
 The training parameters are set according to the parameters specified in the original paper, respectively.
-We use three evaluation metrics, peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and root mean squared error (RMSE), to quantitatively evaluate the generated results to verify the effectiveness of our method.
+We use two evaluation metrics, peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively evaluate the generated results to verify the effectiveness of our method.
 5 Experimental Results
 5.1 Qualitative Results
+Table 1: Quantitative comparison of different approaches on BraTs and ISLSE.
+The values in the table represent the average values for the entire test dataset, and higher values are better.
+Bold indicates the best result, and underlined indicates the suboptimal result.
+BraTs (2023) ISLES (2015)
+Dataset T1–>T2 T1–>FLAIR FLAIR–>DWI FLAIR–>T1
+Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+Ea-GAN 22.43 0.825 21.16 0.773 24.18 0.850 21.35 0.810
+RevGAN 22.61 0.816 21.83 0.769 25.40 0.862 23.64 0.844
+MaskGAN 21.40 0.809 20.41 0.735 25.24 0.856 23.15 0.834
+ResViT 22.89 0.822 21.87 0.786 25.67 0.872 24.90 0.873
+SynDiff 21.68 0.814 20.41 0.755 25.39 0.872 24.32 0.867
+DFBK (ours) 23.61 0.846 22.38 0.797 25.76 0.879 25.42 0.879
 Fig.
@@
 3 (e)).
-ResShift generates a generally correct anatomical structure, but the inconsistency of the brightness of certain areas with the target modality is still a noticeable problem (Fig.
-3 (f)).
 In contrast, the images generated by our proposed method demonstrate significantly improved anatomical accuracy and more precise contrast representation.
-As shown in Figure 3 (g), our approach effectively preserves the structural integrity of anatomical features while maintaining the correct brightness distribution according to target modality.
+As shown in Figure 3 (f), our approach effectively preserves the structural integrity of anatomical features while maintaining the correct brightness distribution according to target modality.
 Furthermore, our method achieves the better performance across the quantitative evaluation metrics, underscoring its superiority in medical image translation.
-Table 2: Quantitative comparison of different approaches on SynthRAD.
-The values in the table represent the average values for the entire test dataset.
-Bold represents the best result and underlined represents the suboptimal result.
-Dataset SynthRAD (2023)
-MR–>CT CT–>MR
-Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
-Ea-GAN 24.75 0.852 0.059 20.39 0.679 0.098
-RevGAN 26.24 0.872 0.051 22.94 0.798 0.073
-MaskGAN 25.56 0.871 0.055 22.04 0.761 0.081
-ResViT 26.92 0.887 0.047 22.88 0.801 0.075
-SynDiff 26.05 0.877 0.052 20.73 0.705 0.094
-ResShift 27.31 0.896 0.045 23.35 0.824 0.071
-DFBK (ours) 27.51 0.897 0.044 23.54 0.828 0.069
-Table 3: Ablation study on Dynamic Frequency Balance (DFB) and Knowledge Guidance (KG). “None” indicates the performances in the absence of DFB and KG.
-Task FLAIR–>DWI FLAIR–>T1
-Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
-None 25.61 0.877 0.055 25.15 0.876 0.062
-+DFB 25.70 0.878 0.055 25.21 0.877 0.062
-+KG 25.68 0.877 0.055 25.31 0.878 0.061
-+DFB, KG 25.76 0.879 0.055 25.42 0.879 0.060
 Fig.
@@
 The anatomical structure generated by ResViT (Figure 4 (d)) and SynDiff (Figure 4 (e)) exhibits noticeable incorrect, with significant distortions in shape and structure.
-Additionally, the brightness in certain regions is excessively dark, which fails to align with the target modality.
-Similarly, the anatomical structure generated by ResShift (Figure 4 (f)) suffers from partial distortions, where some regions exhibit structural inconsistencies.
-Moreover, the brightness distribution is uneven, with certain areas deviating from the expected intensity levels of the target modality.
-In contrast, as shown in Figure 4 (g), our method presents more accurate anatomical structures and achieves better metric values.
-Fig.
-5 shows the results of various methods for medical image translation on SynthRAD.
-The task of CT to MR translation in SynthRAD is very difficult.
+In contrast, as shown in Figure 4 (f), our method presents more accurate anatomical structures and achieves better metric values.
+We also shows the results on the difficult task of CT to MR translation on SynthRAD, as shown in Fig.
+5 in Supplementary Material B.
 Since CT images do not provide information about the internal anatomical structure of the brain, most approaches cannot generate correct or relevant target modality anatomical structures.
@@
 5 (e)) localize the lesion area in generated MR images, they do not accurately reproduce the overall internal anatomy and CT images generated by ResViT miss part of the cranial boundaries.
-ResShift produces MR images with generally correct anatomy but still fails to locate the lesion area, and its CT images are also incomplete (Fig.
+In contrast, our method accurately localizes the MR lesion area, generates correct anatomical structures, and produces CT images with correct cranial boundaries (Fig.
 5 (f)).
-In contrast, our method accurately localizes the MR lesion area, generates correct anatomical structures, and produces CT images with correct cranial boundaries (Fig.
-5 (g)).
 The more visualization results are in Supplementary Material B.
+Table 2: Quantitative comparison of different approaches on SynthRAD.
+Dataset SynthRAD (2023)
+MR–>CT CT–>MR
+Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+Ea-GAN 24.75 0.852 20.39 0.679
+RevGAN 26.24 0.872 22.94 0.798
+MaskGAN 25.56 0.871 22.04 0.761
+ResViT 26.92 0.887 22.88 0.801
+SynDiff 26.05 0.877 20.73 0.705
+DFBK (ours) 27.51 0.897 23.54 0.828
+Table 3: Ablation study on Dynamic Frequency Balance (DFB) and Knowledge Guidance (KG). “None” indicates the performances in the absence of DFB and KG.
+Task FLAIR–>DWI FLAIR–>T1
+Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+None 25.61 0.877 25.15 0.876
++DFB 25.70 0.878 25.21 0.877
++KG 25.68 0.877 25.31 0.878
++DFB, KG 25.76 0.879 25.42 0.879
 5.2 Quantitative Results
 Tab.
-1 shows the mean values for each quantitative metric, where higher PSNR and SSIM values perform better, while lower RMSE value performs better.
+1 and Tab.
+3 shows the mean values for each quantitative metric, where higher PSNR and SSIM values perform better.
 Our method outperforms all other approaches in six translation tasks on three datasets.
@@
 SynDiff, which incorporates GAN principles into a diffusion model, also shows unsatisfactory performance because its more complex training process is more sensitive to data quality.
-ResViT achieves relatively good results by leveraging large model fine-tuning and incorporating prior knowledge for cross-domain image synthesis, and ResShift shows suboptimal performance on most metrics through its efficient diffusion process, but the images generated by ResShift still show brightness inconsistency and anatomical distortion due to insufficient low-frequency information extraction.
+ResViT achieves relatively good results by leveraging large model fine-tuning, but still causes anatomical distortion.
 In contrast, our method effectively achieves frequency balance and has better performance.
@@
 3, we can see that both dynamic frequency balance and knowledge guidance play an important role in generating the correct anatomical structure and affect most indicators, such as PSNR and SSIM.
-At the same time, it can be seen that when the source modality and target modality are significantly different, such as the conversion from FLAIR to T1, the knowledge-guided mechanism plays an important role and can effectively guide the model to generate the correct anatomical structure.
-Table 4: Ablation study on the components in the Dynamic Frequency Balance module. “None” indicates the performances when there are no additional operations in this module.
-Task FLAIR–>DWI FLAIR–>T1
-Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
-None 25.61 0.877 0.055 25.15 0.876 0.062
-+L 25.65 0.878 0.055 25.17 0.876 0,063
-+H 25.67 0.877 0.055 25.12 0.874 0,063
-+NormAttn 25.59 0.878 0.055 25.09 0.874 0.063
-+L, H 25.70 0.878 0.055 25.21 0.877 0.062
-Table 5: Ablation study on the components in the Knowledge-Guided mechanism. “None” indicates the performances when there are no additional operations in this mechanism.
-Task FLAIR–>DWI FLAIR–>T1
-Method PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ RMSE $\downarrow$
-None 25.61 0.877 0.055 25.15 0.876 0.062
-+L 25.66 0.876 0.055 25.21 0.878 0,062
-+H 25.59 0.875 0.055 25.15 0.874 0.062
-+CrossAttn 25.64 0.877 0.055 25.21 0.875 0.062
-+L, H 25.68 0.877 0.055 25.31 0.878 0.061
-The components in DFB. In this section, we will verify the impact of low-frequency and high-frequency modules on model performance respectively.
+At the same time, it can be seen that when the source modality and target modality are significantly different, such as the conversion from FLAIR to T1, the knowledge-guided mechanism plays an important role and can effectively guide the model to generate correct anatomical structures.
+The components in DFB.
 It can be seen from Tab.
@@
 In contrast, the dynamic frequency balance module can decompose and process frequency components more carefully.
-The components in KG. In order to verify the impact of each component of the knowledge-guided mechanism, we conducted an ablation experiment.
+The components in KG.
 As shown in Tab.
@@
 In contrast, the knowledge-guided mechanism processes high and low frequencies separately, effectively improving model performance.
+Table 4: Ablation study on the components in the Dynamic Frequency Balance module. “None” indicates the performances when there are no additional operations in this module.
+Task FLAIR–>DWI FLAIR–>T1
+Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+None 25.61 0.877 25.15 0.876
++L 25.65 0.878 25.17 0.876
++H 25.67 0.877 25.12 0.874
++NormAttn 25.59 0.878 25.09 0.874
++L, H 25.70 0.878 25.21 0.877
+Table 5: Ablation study on the components in the Knowledge-Guided mechanism. “None” indicates the performances when there are no additional operations in this mechanism.
+Task FLAIR–>DWI FLAIR–>T1
+Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+None 25.61 0.877 25.15 0.876
++L 25.66 0.876 25.21 0.878
++H 25.59 0.875 25.15 0.874
++CrossAttn 25.64 0.877 25.21 0.875
++L, H 25.68 0.877 25.31 0.878
 6 Conclusion
@@
 As shown in Fig.
-6, 7 and 8, our method consistently outperforms other methods for medical image translation on BraTs, ISLES, and SynthRAD datasets, demonstrating superior anatomical accuracy, and better quantitative metrics.
+5, 6, 7 and 8, our method consistently outperforms other methods for medical image translation on BraTs, ISLES, and SynthRAD datasets, demonstrating superior anatomical accuracy, and better quantitative metrics.
 Unlike other methods, our approach effectively maintains structural integrity, accurately locates lesion areas, and ensures realistic conversion between imaging modalities.
 Therefore, our method is robust and reliable in medical image translation and has great clinical application value.
-Figure 6: Additional results on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours).
+Figure 5: Comparison of approaches on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours).
+The first and third rows show the results of CT to MR and MR to CT translation, respectively.
+The first column is the source image, and the second column is the target ground-truth.
+The second and fourth rows show the zoomed-in anatomical structures for the first and third rows, respectively.
+Figure 6: Additional results on BraTs: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours).
 The first and third rows show the results of T1 to T2 and T1 to FLAIR translation, respectively.
@@
 The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
-Figure 7: Additional results on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours).
+Figure 7: Additional results on ISLES: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours).
 The first and third rows show the results of FLAIR to DWI and FLAIR to T1 translation, respectively.
@@
 The color indicates the degree of difference from small to large, with the brighter color (e.g., red) reflecting the larger difference and vice versa.
-Figure 8: Additional results on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff, (f) ResShift, (g) DFBK (ours).
+Figure 8: Additional results on SynthRAD: (a) Ea-GAN, (b) RevGAN, (c) MaskGAN, (d) ResViT, (e) SynDiff (f) DFBK (ours).
 The first and third rows show the results of CT to MR and MR to CT translation, respectively.